# Stylometry recognizes human and LLM-generated texts in short samples

### v1

@@ 2.4 LLMs safety and ethics
Another survey – on Large Language Model (LLM) Security and Privacy: The Good, the Bad, and the Ugly (Yao et al., 2024) offers a detailed exploration of the security and privacy dimensions associated with LLMs. It assesses how LLMs can both enhance and threaten cybersecurity in various applications. The authors categorize their findings into beneficial uses ("The Good"), such as improving code security and data privacy, offensive applications ("The Bad"), like their use in user-level attacks due to their sophisticated reasoning capabilities, and inherent vulnerabilities ("The Ugly") that could be exploited maliciously. The survey emphasizes the dual nature of LLMs in cybersecurity, showcasing their potential to advance security measures while also posing significant risks if not carefully managed and regulated. Furthermore, it identifies areas needing further research, such as model and parameter extraction attacks and the development of safe instruction tuning, underlining the complexity and evolving nature of LLM applications in security contexts.
@@
Last, but not least – ChatGPT and the rise of large language models: the new AI-driven infodemic threat in public health by De Angelis et al. (2023) examines the dual-edged impact of LLMs on public health. It acknowledges the potential of LLMs to aid scientific research through their ability to process and generate large amounts of data quickly. However, it critically highlights the risk of an "AI-driven infodemic," where the rapid and widespread dissemination of misinformation could be facilitated by these same technologies. The paper calls for urgent policy actions to mitigate these risks, emphasizing the need for a balanced approach in harnessing the benefits of LLMs while safeguarding against their potential to undermine public health and the integrity of scientific research. This includes the establishment of regulatory frameworks and the proactive monitoring of the use of LLMs to prevent the spread of false information.
@@ 3 Metholodogy
The process of the proposed solution is divided into several steps. The data acquisition is explained in the first part of the chapter (3.1). The data was next cleaned up and extended by the summaries generated with various text summarization methods (3.5). In the next step, we added additional short terms descriptions generated using different language models (3.2). Finally, based on the stylometric features, we differentiate between the texts generated by the models and the humans (3.3, 3.4).
@@ 3.1 Dataset
The dataset is based on the Wikipedia terms using two different Python libraries: datasets from HuggingFace [^1] https://huggingface.co (Lhoest et al., 2021) and Wikipedia-API. We obtained 1500 terms using the first method and 1048 terms using the second. In the first method we used the dataset from 2022, it is named 20220301.simple. The final dataset used in this paper consists of 2439 terms. The number is a result of the preprocessing part and the removal of all examples that did not meet one of the following requirements:
@@
    the first 10 sentences does not include references (bibliography).
@@
Each term description that did not fulfill the above requirements was removed from the dataset. Before the above validation, non-latin letters were removed, characters like duplicated whitespaces were removed, this include brackets, semicolons, and dots.
3.2 Language Models
@@
We choose a few language models, including the open and API-based ones. We used the ChatGPT API for two models: GPT-3.5-turbo, and GPT-4 (Liu et al., 2023). LLaMa 2 and 3 with 7 and 8 billion parameters, respectively (Touvron et al., 2023). In this case, we used the Ollama [^2] https://ollama.com/ library. For the other two models: Orca (Mukherjee et al., 2023) and Falcon (Almazrouei et al., 2023) we used the GPT4All library (Anand et al., 2023). The models we used have 8 and 11 billion parameters respectively.
@@
We used two prompts that were sent to each of the models. The first one is a simple ask for term explanation in 10 sentences. The exact prompt is the following: Please describe i n 10 sentences as plain text what <term> is. The second prompt is a request for a text similar to the Wikipedia page. The exact prompt is the following: Please describe as it would be the Wikipedia page in 10 sentences what <term> is. The reason of having two prompts is that the term explanation can be potentially easier to be recognized when compared with a model generated text. That is why the Wikipedia page-like response is compared.
@@ 3.3.1 StyloMetrix
StyloMetrix is an open-source stylometric text analysis library. Covers various grammatical, syntactic, and lexical aspects. StyloMetrix allows allowing feature engineering and interpretability. Stylometry involves the analysis of linguistic features to characterize the style of texts. Previous tools like ’stylo’ package in R (Eder et al., 2016) provide quantitative text analysis but lack certain metrics and usability features that StyloMetrix offers. It is based on the spaCy model for English and generates normalized vectors for input texts, allowing comparison across texts of different lengths and genres. Vectors are designed to be interpretable at different levels. Metrics that are available for the English language:
@@ 3.3.2 CLARIN-PL’s stylometric pipeline
We used a modular Python pipeline for interpretable stylometric analysis developed for CLARIN-PL [^3] https://gitlab.clarin-pl.eu/stylometry/cl_explainable_stylo(Ochab & Walkowiak, 2024). The pipeline connects text preprocessing and linguistic feature extraction with various NLP tools, classifiers, an explainability module, and visualization. At present, we use spaCy (Montani et al., 2023) model ’en_core_web_lg’ for preprocessing steps (including tokenisation, named entity recognition, dependency parsing, part-of-speech and morphology annotation), Light Gradient-Boosting Machine (LGBM) (Ke et al., 2017) as the state-of-the-art boosted trees classifier, Shapley Additive Explanations (SHAP) (Lundberg et al., 2020) for computing explanations, and Scikit-learn (Pedregosa et al., 2011) for feature counting and cross-validation. The visualisation functions, showing general and detailed explanations of what linguistic features make texts differ, utilise spaCy and SHAP.
@@ 3.4 Classification
The first method chosen is a simple decision tree classifier from the popular Python sklearn [^4] https://scikit-learn.org library. It was used with the default parameters such as the Gini impurity method, the minimum samples in the split set to 2, and the split strategy set to best. The test and train sets that was used in a split of 70% to 30% with a 10 cross validation.
The LGBM classifier was used with the following settings: DART boosting, maximal depth of the tree model ("max_depth" = 5), maximal number of leaves per tree ("num_leaves" = 5), default number of boosting iterations, increased "learning_rate" = 0.5, enabled bagging (randomly selecting part of data without resampling with "bagging_freq" = 3 and "bagging_fraction" = 0.8), and number of classes in the multiclass scenario ("num_class" = 7).
We used the group cross-validation (CV) scheme by using 10-fold CV for test error estimation. Group CV makes sure that a given topic of the summary never appears both in the train and test set. The reported scores are averages over the CV loop. Training and test set sizes in each fold were $43904390$ and $488488$ samples for binary classification and, respectively, $1536515365$ and $17081708$ for multiclass classification.
@@
3.5 Text Summarizers
We have used four text summarization methods for comparison reason. This includes a very popular Python method in the gensim library (Řehůřek & Sojka, 2010). It is already outdatted as there are more complex methods based on transformers that are supposed to give better results. This method is the T5 (Raffel et al., 2020) and BART summarizers (Lewis et al., 2019). Both are used in our research. The last summarization method is called sumy and is implemented in the sumy [^5] https://pypi.org/project/sumy library.
Every summarization method is fed with the Wikipedia terms descriptions, but each summarization method does have different parameters to be set. We tried to set such parameters to get a summary of about 10 sentences for each term. The gensim summarizer does have a number of sentences parameter, but we did not set it to an exact number. It produced a sufficient number of sentences and in case if it exceeded we just drop the excess sentences. For the T5 and BART summarizers we got the best results with setting the maximum number of characters to 1000.The lenght penality parameter and number of beans were left to the standard values of 2.0 and 4 respectively. Sumy does have a parameter that allows one to set the exact number of sentences. We set it to 10.
wiki gpt3.5 gpt4 llama2 llama3 orca falcon
@@ 3.5 Text Summarizers
Table 1: Accuracy of decision tree classification of two type of texts generated using different prompts
@@ 4 Results
4.1 Decision trees binary classification
@@
Decision trees are known to be used for measure features importances. In our first experiment the most significant stylometric features are as following:
@@
These four features were used for the 1vs1 classification. The worst results were achieved for the second prompt with the following comparisons: orca vs. llama2, llama3 vs. gpt4, falcon vs. llama2, and falcon vs. orca. In the first two cases the results were about 52% and 56% accordingly. We can come to a conclusion that in both cases the recognition is very limited or even does not recognize. Majority of model 1vs1 recognitions are between 70% and 85%. The best results are the llama2 vs. gpt3.5, and orca vs. gpt3.5 for both prompt. The accuracy is about 92% for the first prompt, and about 89% for the second prompt. What is worth attention are the results in recognition of models’ generated text and the Wikipedia text where the lowest accuracy is about 73%, but the majority is above 85%, with best results achieved for orca and llama2, 95% and 96% accrodingly.
@@ 4.2.1 StyloMetrix features
wiki gpt3.5 gpt4 llama2 llama3 orca falcon
@@ 4.2.2 Frequency-based features
Table 3 shows the normalised confusion matrix. Again, Wikipedia has the highest accuracy and the LLM most often misclassified as it is GPT-4. The most often confused pairs of models are Falcon and Orca, GPT-3 and GPT-4, Llama3 and GPT-3.
4.3 Binary classification with LGBM
4.3.1 StyloMetrix features
Table 4 shows CV-averaged accuracy between all pairs of classes. The LLM most often misclassified as the real Wikipedia are GPT-4 and Llama3 (cf. Tables 3-4). Llama2 and Orca were the hardest to distinguish. GPT models and Llama3, as well as Orca and Falcon are also confused often.
wiki gpt3.5 gpt4 llama2 llama3 orca falcon
Stylometrix features
wiki 0.97 0.94 0.99 0.95 0.99 0.98
gpt3.5 0.87 0.99 0.88 0.99 0.98
gpt4 0.99 0.85 0.99 0.98
llama2 0.99 0.77 0.90
llama3 0.99 0.98
orca 0.87
falcon
Frequency-based features
wiki 0.99 0.98 1.00 0.99 1.00 1.00
gpt3.5 0.90 0.98 0.91 0.98 0.97
gpt4 0.99 0.93 0.99 0.98
llama2 0.99 0.79 0.84
llama3 1.00 0.99
orca 0.86
falcon
Table 4: Accuracy between pairs of models of binary LGBM classifier using StyloMetrix features. Average over 10 CV folds.
4.3.2 Frequency-based features
Table 4 shows accuracy between all pairs of classes. LLMs are hardly confused with the real Wikipedia at all. As before, the most often confused pairs of models were GPT models and Llama3, as well as the triplet Llama2, Orca and Falcon.

### v2

@@ 2.7 LLMs safety and ethics
The application of stylometry to LLMs is particularly important given the potential risks associated with their misuse, such as the generation of misleading information, deepfake text, or malicious content, as described below. We note, however, that mere detection that a text has been machine-generated – which is the objective of the present paper – does not imply that it is untrustworthy or malicious. Schuster et al. (2020) reported that, even though human language tends to stylistically change when deceiving, stylometry fails to detect malicious use of (now perhaps obsolete) LLMs, and that such issues involve a whole ecosystem of fraud (including among others fact-checking, users’ feedback, and content propagation through social networks).
@@
Another survey – on Large Language Model (LLM) Security and Privacy: The Good, the Bad, and the Ugly (Yao et al., 2024) offers a detailed exploration of the security and privacy dimensions associated with LLMs. It assesses how LLMs can both enhance and threaten cybersecurity in various applications. The authors categorize their findings into beneficial uses ("The Good"), such as improving code security and data privacy, offensive applications ("The Bad"), like their use in user-level attacks due to their sophisticated reasoning capabilities, and inherent vulnerabilities ("The Ugly") that could be exploited maliciously. The survey emphasises the dual nature of LLMs in cybersecurity, showcasing their potential to advance security measures while also posing significant risks if not carefully managed and regulated. Furthermore, it identifies areas needing further research, such as model and parameter extraction attacks and the development of safe instruction tuning, underlining the complexity and evolving nature of LLM applications in security contexts.
@@
Last, but not least – ChatGPT and the rise of large language models: the new AI-driven infodemic threat in public health by De Angelis et al. (2023) examines the dual-edged impact of LLMs on public health. It acknowledges the potential of LLMs to aid scientific research through their ability to process and generate large amounts of data quickly. However, it critically highlights the risk of an “AI-driven infodemic”, where the rapid and widespread dissemination of misinformation could be facilitated by these same technologies. The paper calls for urgent policy actions to mitigate these risks, emphasizing the need for a balanced approach in harnessing the benefits of LLMs while safeguarding against their potential to undermine public health and the integrity of scientific research. This includes the establishment of regulatory frameworks and the proactive monitoring of the use of LLMs to prevent the spread of false information.
@@ 3 Metholodogy
The process of the proposed solution is divided into several steps. The data acquisition and cleaning is explained in the first part of the chapter (3.1.1). The data was next extended by the summaries generated with various text summarisation methods (3.1.2). In the next step, we added additional short terms descriptions generated using different language models (3.1.3). Finally, based on the stylometric features, we differentiate between the texts generated by the models and the humans (3.3, 3.4).
@@ 3.1 Dataset
3.1.1 Human texts: Wikipedia summaries
The dataset is based on Wikipedia terms using two different Python libraries: datasets from HuggingFace [^1] https://huggingface.co (Lhoest et al., 2021) and Wikipedia-API. In the first method we used the dataset from 2022 named 20220301.simple. Similarly to how Bevendorff et al. (2024) collected their data, our choice was dictated by the date of GPT-3.5 release, so that we avoid contaminating the human-authored and edited texts with MGT in the view of its increased presence in Wikipedia (Brooks et al., 2024). We obtained 1500 terms using the first method and 1048 terms using the second. The final dataset used in this paper consists of 2439 terms. The number is a result of the preprocessing part and the removal of all examples that did not meet one of the following requirements:
@@ 3.1.1 Human texts: Wikipedia summaries
    the first 10 sentences do not include references (bibliography).
@@
Each term description that did not fulfil the above requirements was removed from the dataset. Before the above validation, non-latin letters were removed, and characters like duplicated whitespaces were removed, including brackets, semicolons, and dots. For classification purposes, the texts were shortened to a maximum of 18 sentences (this is the maximum number of sentences generated by GPT models, despite the 10-sentence limit). Additionally, we removed several outlying texts (which had high sentence counts, mainly due to improper spaCy segmentation of lists or enumerations), resulting in 2424 terms.
3.1.2 Text summarizers
We used four text summarisation methods for comparison: (1) A very popular Python method in the gensim library (Řehůřek & Sojka, 2010). It is already outdated, as there are more complex methods based on transformers that reportedly give better results. (2-3) Transformer-based T5 (Raffel et al., 2020) and BART summarizers (Lewis et al., 2019). (4) The last summarization method is called sumy and is implemented in the sumy [^2] https://pypi.org/project/sumy library.
Every summarisation method is provided with the Wikipedia terms descriptions, but each has different parameters to be set. We tried to set such parameters to obtain a summary of about 10 sentences for each term. The gensim summarizer does have a ‘number of sentences’ parameter, but we did not set it to an exact number. It produced a sufficient number of sentences and in case it exceeded the limit, we just dropped the excess sentences. For the T5 and BART summarizes we got the best results with setting the maximum number of characters to 1000. The length penalty parameter and number of beans were left to the standard values of 2.0 and 4, respectively. Sumy has a parameter that allows one to set the exact number of sentences, which we set to 10.
Table 1 shows basic statistics of the dataset. In particular, T5 tends to produce a highly varying summary length, both in terms of tokens and sentences, and amount of punctuation. The reason is its failure, resulting in repetition of the same letters or words and a number of full stops. The other summarisers do not produce such artefacts, with BART generating a low number of short sentences, Gensim a low number of relatively longer sentences, and Sumy a larger number of sentences.
Gensim Sumy T5 BART
#tokens 71 ± 36 249 ± 62 220 ± 110 61 ± 12
fraction of punctuation [%] 13.4 ± 4.9 13.8 ± 4.0 28 ± 11 13.1 ± 4.3
#tokens / sentence 28.8 ± 9.8 24.4 ± 6.0 14.0 ± 13.0 19.7 ± 5.5
#sentences 2.5 ± 1.0 10.4 ± 2.2 21.0 ± 15.0 3.19 ± 0.7
max. #sentences 9 41 138 7
Table 1: Basic summariser dataset statistics: total number of tokens (including punctuation), fraction of punctuation tokens in the total token count, mean number of tokens in a sentence, number of sentences, and the maximal number of sentences. The numbers are averages and standard deviations across all documents.
3.1.3 LLM-generated descriptions
@@
Large language models were used to generate term descriptions from scratch, i.e., they were provided only with terms they were prompted to describe, but not with any part of the Wikipedia articles. We chose six language models, including the open and API-based ones. We used the ChatGPT API for two models: GPT-3.5-turbo, and GPT-4 (Liu et al., 2023a). LLaMa 2 and 3 with 7 and 8 billion parameters, respectively (Touvron et al., 2023). In this case, we used the Ollama [^3] https://ollama.com/ library. For the other two models: Orca (Mukherjee et al., 2023) and Falcon (Almazrouei et al., 2023) we used the GPT4All library (Anand et al., 2023). The models we used had 8 and 11 billion parameters, respectively. We used the GPT4All library to execute LLaMa2, LLaMa3, Orca, and Falcon. We used the default temperature value. Based on the documentation, the temperature was set to 0.7. For GPT3.5 and 4, the temperature setting was 0.7 (currently, the default value for the API for GPT4o and newer is set to 1; OpenAI, 2025).
@@
We used two prompts that were sent to each of the models. The first one is a simple ask for a term explanation in 10 sentences. The exact prompt is the following: Please describe in 10 sentences as plain text what <term> is. The second prompt is a request for a text similar to the Wikipedia page. The exact prompt is the following: Please describe as it would be the Wikipedia page in 10 sentences what <term> is. The reason for having two prompts is that the term explanation can be potentially easier to be recognized when compared with a model-generated text. That is why the Wikipedia page-like response is compared.
Table 2 shows basic statistics of the dataset. In particular, the GPT models and LLaMa 3 kept very close to the 10-sentence limit, while the other models tended to produce shorter sentences and paragraphs.
wiki GPT-3.5 GPT-4 LLaMa 2 LLaMa 3 Orca Falcon
#tokens 243 ± 56 223 ± 35 218 ± 27 152 ± 17 249 ± 35 144 ± 27 152 ± 39
fraction of punctuation [%] 13.5 ± 3.3 11 ± 2.3 12.2 ± 2.3 11.3 ± 2.9 11.8 ± 2.3 12.3 ± 3.4 10.7 ± 3.1
#tokens / sentence 24.0 ± 5.5 23.1 ± 2.8 22.3 ± 2.6 22.0 ± 3.5 24.8 ± 3.1 22.3 ± 3.9 22.3 ± 3.5
#sentences 10.19 ± 0.94 9.7 ± 1.2 9.78 ± 0.79 7.0 ± 1.3 10.04 ± 0.82 6.6 ± 1.7 7.0 ± 2.0
max. #sentences 18 18 14 16 17 16 17
Table 2: Basic LLM dataset statistics: total number of tokens (including punctuation), fraction of punctuation tokens in the total token count, mean number of tokens in a sentence, number of sentences, and the maximal number of sentences. The numbers are averages and standard deviations across all documents.
3.2 Benchmarks
As external benchmark for machine-generated text detection we have used AuTexTification Sarvazyan et al., 2023a. AuTexTification contains two shared tasks in two languages (English and Spanish): (i) MGT detection (a binary classification of texts written by human and a language model) and (ii) MGT attribution (classification of six models). Importantly, the first task uses a balanced multi-domain (tweets, how-to articles, legal documents, reviews and news) and multi-model (BLOOM: BLOOM-1B7, BLOOM-3B, BLOOM-7B1, and GPT-3: babbage, curie, and text-davinci-003) corpus, where only the first three domains appear in the training data and the last two in the test data.
Benchmark Human LLMs LLMs Type Language Domain
AuTexTification 1 $\sim$28k $\sim$28k BLOOM-1B1, BLOOM-3B, BLOOM-7B1, Babbage, Curie, text-davinci-003 English tweets, how-to, news, legal, reviews
Table 3: Benchmark used.
@@ 3.3.1 StyloMetrix
StyloMetrix is an open-source stylometric text analysis library. Covers various grammatical, syntactic, and lexical aspects. StyloMetrix allows allowing feature engineering and interpretability. Stylometry involves the analysis of linguistic features to characterize the style of texts. Previous tools like ‘stylo’ package in R (Eder et al., 2016) provide quantitative text analysis but lack certain metrics and usability features that StyloMetrix offers. It is based on the spaCy model for English and generates normalized vectors for input texts, allowing comparison across texts of different lengths and genres. Vectors are designed to be interpretable at different levels. Metrics that are available for the English language:
@@ 3.3.2 CLARIN-PL’s stylometric pipeline
We used a modular Python pipeline for interpretable stylometric analysis developed for CLARIN-PL [^4] https://gitlab.clarin-pl.eu/stylometry/cl_explainable_stylo(Ochab & Walkowiak, 2024). The pipeline connects text preprocessing and linguistic feature extraction with various NLP tools, classifiers, an explainability module, and visualization. At present, we use spaCy (Montani et al., 2023) model ‘en_core_web_lg’ for preprocessing steps (including tokenisation, named entity recognition, dependency parsing, part-of-speech and morphology annotation), Light Gradient-Boosting Machine (LGBM) (Ke et al., 2017) as the state-of-the-art boosted trees classifier, Shapley Additive Explanations (SHAP) (Lundberg et al., 2020) for computing explanations, and Scikit-learn (Pedregosa et al., 2011) for feature counting and cross-validation. The visualisation functions, showing general and detailed explanations of what linguistic features make texts differ, utilise spaCy and SHAP.
@@ 3.4 Classification
The first method chosen is a simple decision tree classifier from the popular Python sklearn [^5] https://scikit-learn.org library. It was used with the default parameters such as the Gini impurity method, the minimum samples in the split set to 2, and the split strategy set to best. The test and train sets were used in a split of 70% to 30% with a 10-fold cross-validation (CV).
The LGBM classifier was used with the following settings: DART boosting, maximal depth of the tree model ("max_depth" = 5), maximal number of leaves per tree ("num_leaves" = 5), default number of boosting iterations, increased "learning_rate" = 0.5, enabled bagging (randomly selecting part of data without resampling with "bagging_freq" = 3 and "bagging_fraction" = 0.8), and number of classes in the multiclass scenario ("num_class" = 7). Further hyperparameter optimisation is possible, but was not performed in this study.
We used the group cross-validation scheme by using 10-fold CV for test error estimation. Group CV makes sure that a given topic of the summary never appears both in the train and test set. The reported scores are averages over the CV loop. Training and test set sizes in each fold were $4390$ and $488$ samples for binary classification and, respectively, $15365$ and $1708$ for multiclass classification.
@@
wiki GPT-3.5 GPT-4 LLaMa 2 LLaMa 3 Orca Falcon
@@
Table 4: Accuracy of binary text classification with decision trees. Each table entry corresponds to a task, where class 1 and 2 are column and row model labels, respectively. Texts generated by different prompts are analysed separately.
@@ 4 Results
4.1 Binary classification with decision trees
@@
Decision trees are known to be used to measure feature importance. In our first experiment the most significant stylometric features are as follows:
@@
These four features were used for the binary classifications. The worst results were achieved for the second prompt with the following pair of classes: Orca and LLaMa 2, LLaMa 3 and GPT-4, Falcon and LLaMa 2, and Falcon and Orca. In the first two cases the results were about 52% and 56% accordingly. We can conclude that in both cases the recognition is very limited or even fails. Majority of model binary recognitions are between 70% and 85%. The best results are for distinguishing LLaMa 2 from GPT-3.5, and Orca from GPT-3.5 for both prompts. The accuracy is about 92% for the first prompt, and about 89% for the second prompt. What is worth attention are the results in recognition of models’ generated text and the Wikipedia text where the lowest accuracy is about 73%, but the majority is above 85%, with best results achieved for Orca and LLaMa 2, 95% and 96% accordingly.
4.2 Binary classification with LGBM
4.2.1 StyloMetrix features
Table 5 shows CV-averaged accuracy between all pairs of classes. The LLM most often misclassified as the real Wikipedia are GPT-4 and LLaMa 3 (cf. Tables 7-5). LLaMa 2 and Orca were the hardest to distinguish. GPT models and LLaMa 3, as well as Orca and Falcon are also confused often.
wiki GPT-3.5 GPT-4 LLaMa 2 LLaMa 3 Orca Falcon
Stylometrix features
wiki 0.97 0.94 0.99 0.95 0.99 0.98
GPT-3.5 0.87 0.99 0.88 0.99 0.98
GPT-4 0.99 0.85 0.99 0.98
LLaMa 2 0.99 0.77 0.90
LLaMa 3 0.99 0.98
Orca 0.87
Falcon
Frequency-based features
wiki 0.99 0.98 1.00 0.99 1.00 1.00
GPT-3.5 0.90 0.98 0.91 0.98 0.97
GPT-4 0.99 0.93 0.99 0.98
LLaMa 2 0.99 0.79 0.84
LLaMa 3 1.00 0.99
Orca 0.86
Falcon
Table 5: Accuracy of binary text classification with LGBM using StyloMetrix features. Each table entry corresponds to a task, where class 1 and 2 are column and row model labels, respectively. The results are averages over 10 CV folds.
4.2.2 Frequency-based features
Table 5 shows the accuracy between all pairs of classes. LLMs are hardly confused with the real Wikipedia at all. As before, the most often confused pairs of models were GPT models and LLaMa 3, as well as the triplet LLaMa 2, Orca, and Falcon.
@@ 4.3.1 StyloMetrix features
wiki GPT-3.5 GPT-4 LLaMa 2 LLaMa 3 Orca Falcon
@@ 4.3.2 Frequency-based features
Table 7 shows the normalised confusion matrix. Again, Wikipedia has the highest accuracy and the LLM most often misclassified as it is GPT-4. The most often confused pairs of models are Falcon and Orca, GPT-3.5 and GPT-4, LLaMa 3 and GPT-3.5.
4.4 Robustness testing
For brevity, we provide robustness testing only for the LGBM model with frequency-based features, and only for the binary detection of human- and machine-generated texts. Since the test sets contains only one class (machine-generated texts), we provide the value of recall of that class and the validation recall for comparison, see Table 8.
4.4.1 Testing on unseen models
The test assumes that in training the model can only access data on man-made texts and on five out of six LLMs. The features are chosen and fixed at this stage and the training recall is computed. Testing is performed on the single LLM previously unseen by the model. The cross-validation here regards the training only, i.e., for each unseen LLM there were 10 classifiers trained on subsets of the training set (and evaluated on the validation set as shown in Table 8), while the test set remained the same. The standard deviations of recall are computed over these 10 folds.
The largest drop in performance can be seen for GPT-4 and LLaMa 3 models.
Recall [%] Validation Test DIPPER Parrot
GPT-3.5 99.11 ± .36 99.6 ± .17 99.95±.062 99.971 ± 0.044
GPT-4 99.49 ± .24 88.2 ± 1.2 99.95±.045 98.81 ± 0.14
LLaMa 2 99.17 ± .39 99.61 ± .11 99.922±.065 99.992 ± 0.017
LLaMa 3 99.24 ± .24 94.13 ± .72 99.9±.064 99.736 ± 0.098
Orca 99.18 ± .32 99.79 ± .16 99.87±.17 99.996 ± 0.013
Falcon 99.14 ± .30 99.691 ± .056 99.81±.11 99.955 ± 0.03
Table 8: Recall on (i) unseen LLMs and on texts paraphrased with (ii) DIPPER and (iii) Parrot. (i) Unseen LLM test set contained only the given model, while training set contained the other models and the human Wiki summaries. (ii)-(iii) Paraphrased test set contained only the paraphrases, while the training set contained all the models and human texts. The values are mean and standard deviation over CV folds.
4.4.2 Testing on paraphrased texts
Following Sadasivan et al. (2025) we performed a paraphrase attack using DIPPER Krishna et al. (2023), a 11B paraphrasing model, and Parrot Damodaran (2021), a T5-based paraphrase model. Reportedly, in a small sample, DIPPER had shown in human evaluation that the content was satisfactorily preserved in about 70% of the samples and grammar quality was satisfactory in 88%. We used no recursive paraphrasing.
The test assumes that, in training, the model can access all unparaphrased data (human and all six LLM-generated texts). The features are chosen and fixed at this stage. Testing is performed on all paraphrases. As above, the cross-validation here regards the training only. The results are shown in Table 8.
In general, paraphrasing resulted in a higher detection rate than for unparaphrased texts. The only exception is Parrot paraphrasing GPT-4, where a less than 1% drop in recall occurred.
4.4.3 Testing on cross-domain benchmark
The results obtained on the AuTexTification task are shown in Table 9. Our classifiers used only the training data provided in the task description according to the task constraints.
Classifier F1
Top 0.81
LR 0.66
Mikros et al. (2023) 0.61
StyloMetrix 0.48
Frequencies 0.54
Table 9: AuTexTification benchmark results. Macro-F1 score. For comparison submissions to Sarvazyan et al. (2023a) are presented: the top-ranked, the logistic regression (LR) baseline and results by Mikros et al. (2023) (an ensemble of stylometric features and transformers).